OpenStudy (anonymous):

derivative of (5x^2 -2x)^3

OpenStudy (experimentx):

use chain rule

OpenStudy (hoblos):

(u^n)' = n(u^(n-1))(u') using this ((5x^2 -2x)^3)' = 3((5x^2 -2x)^2)(10x-2)

OpenStudy (saifoo.khan):

2(5x^2 - 2x)^2 (10x - 2)

OpenStudy (saifoo.khan):

Ops, that's a 3.

OpenStudy (anonymous):

\[\LARGE (u^n)'=n \cdot u^{n-1} \cdot u'\]